OpenStudy (anonymous):

10*9*8*7*6*5*4*3*2*1 = 10! Can this be true?! Why or why not?

OpenStudy (anonymous):

yeah it is true

OpenStudy (jamesj):

That's the standard definition of 10! so yes, it's definitely true.

OpenStudy (anonymous):

that's a nice joke. exclamation point mixed for factorial .

OpenStudy (anonymous):

Yes!!! The expression "x!" means x(x-1)(x-2)(x-3)...1 So 10! means 10(10-1)(10-2)(10-3)(10-4)(10-5)(10-6)(10-7)(10-8)(10-9)= 10*9*8*7*6*5*4*3*2*1=10!, which means it is correct.

OpenStudy (anonymous):

When you're in math long enough, you start seeing exclamation points in English papers and wonder why kids are using factorials in research papers.

OpenStudy (anonymous):

haha :P true!!!

OpenStudy (anonymous):

what?! 10*9 is 90 *8 is 720 *7 is 5,040 *6 is 30,240 *5 is 151,200 *4 is 604,800 *3 is 1,814,400*2 is 3,628,800 *1 is 3,628,000. soo, no.

OpenStudy (anonymous):

^lol

OpenStudy (anonymous):

its true

OpenStudy (anonymous):

10!=3628800 and the product of given expresion is also 3628800